OpenStudy (anonymous):

Give the smallest value of x for which the derivative of 3 f (x) = x + 3x + 2 is 0. a. -1 b. 1 c. -2 d. 2 e. 0 f. None of these. Entire class is worried abt this one! Isn't the answer f?

myininaya (myininaya):

is that 3*f(x)?

OpenStudy (anonymous):

sorry!

myininaya (myininaya):

3f(x)=4x+2?

OpenStudy (anonymous):

f(x) = x^3 +3x +2

OpenStudy (anonymous):

great catch!

myininaya (myininaya):

ok f'(x)=3x^2+3 3x^2+3=0 3(x^2+1)=0 x^2+1=0 x=-1 and x=1 smallest is x=-1

OpenStudy (anonymous):

oh I see, what you did! brilliant! makes complete sense, thanks.

myininaya (myininaya):

lol they just want the smallest value satisfying f'(x)=0

OpenStudy (anonymous):

thats; what I did initially. Butm the entire class posted going nuts that it was f blah blah blah.... :-)

myininaya (myininaya):

well gj next time you tell them that there all wrong lol

OpenStudy (anonymous):

@.@ x^2+1=0 =>x is +i,-i isn't it?

myininaya (myininaya):

omg i suck x^2+1=0 x= + or - i

myininaya (myininaya):

i need to go to the bed gj kiwi

myininaya (myininaya):

so the class is right it is f

myininaya (myininaya):

lol

OpenStudy (anonymous):

haha

myininaya (myininaya):

don't tell them they are wrong

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

I need a little review... what's the derivative of a constant? is it 0?

myininaya (myininaya):

yes kiwi

myininaya (myininaya):

derivative=slope slope of a horizontal line (y=c , c is constant) is 0 so the derivative is 0

OpenStudy (anonymous):

thanks